OpenStudy (anonymous):

a/2 - 2a/8 + 6/4 = a + 2/4 Is the answer .25?

OpenStudy (cwrw238):

try plugging in a = 0.25 into the equation and see if it fits

OpenStudy (anonymous):

6/4-2/4=a-a/2+2a/8 1= a( 1-1/2+2/8) a=4/3 right ans